Find a formula for the length of the altitude to the hypotenuse of a right triangle whose legs have lengths of A and B.
Found 2 solutions by rfer, richard1234:
Answer by rfer(16322)

a^2+b^2=c^2
change to
c^2-a^2=b^2

Answer by richard1234(7193)

Let h be the altitude to the hypotenuse. Using the area for a triangle, we have

.

By the Pythagorean theorem, so